# Entropic Risk Optimization in Discounted MDPs: Sample Complexity Bounds with a Generative Model

### v2

@@ Acknowledgments
Appendix A Technical Lemmas
@@
$\log(1-x)$ $\geq-x-x^{2}+x^{3}\quad$ $\forall x\in[0,\frac{1}{5}]$
$\log(1-x)$ $\geq-x-2x\quad$ $\forall x\in[0,\frac{1}{2}]$
$\log(1+x)$ $\geq x-x^{2}\quad$ $\forall x\in[0,\infty)$
$\log(1+x)$ $\geq\frac{x}{2}\quad$ $\forall x\in[0,1]\,.$
@@
We only prove the first claim, as the rest could be proven using the technique after some elementary calculations.
Let f(x)=(1x)f(x)=(1-x) and $g(x)=-x-x^{2}+x^{3}$. It holds that $f(0)=g(0)$, and since we have $f^{\prime}(x)=\frac{1}{1-x}$ and $g^{\prime}(x)=-1-2x+3x^{2}$, it follows easily that
$f^{\prime}(x)\geq g^{\prime}(x)\Leftrightarrow 0\leq x(1-5x+3x^{2})$
@@
Define $f(x)=1-(1-x)^{\alpha}-\frac{x\alpha}{2}$. Since $^{\prime\prime}(x)=-\alpha(\alpha-1)(1-x)^{\alpha-2}<0, ff$ is strictly concave. Further, since $f(0)=0$ and $(\frac{1}{\alpha})=\frac{1}{2}(1-\frac{1}{\alpha})^{\alpha}>\frac{1}{2}-\frac{1}{e}>0, ff$ is positive on the interval $0,\frac{1}{\alpha}]$ and the result follows. ∎
Appendix B Risk Measures
In this section we give a very brief introduction to risk measures with proper definitions and key examples to the extend needed. The reason for this is that the literature is varied and that some inequalities and vocabulary change slightly, in particular over the choice of working with rewards or losses.
Definition 2.
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a background probability space and $\mathcal{M}$ some convex cone of random variables defined on the background space. That is for any $X,Y\in\mathcal{M}$ and $\lambda>0$ it holds that $X+Y\in\mathcal{M}$ and $\lambda X\in\mathcal{M}$. A functional $\psi:\mathcal{M\rightarrow\mathbb{R}}$ is said to be a risk measure if it satisfies the following properties:
$\psi(0)=0\qquad$ (Normalization) (20)
$\text{if }X\leq Y\text{ then }\psi(X)\geq\psi(Y)\qquad$ (Monotonicity) (21)
$\psi(X+c)=\psi(X)-c\quad\forall c\in\mathbb{R}\qquad$ (Translation invariance) (22)
Any risk measure that in addition satisfies the properties
$\psi(cX)=c\psi(X)\quad\forall a>0\qquad$ (Positive homogeneity) (23)
$\psi(X+Y)\leq\psi(X)+\psi(Y)\qquad$ (Subadditivity) (24)
is called a coherent risk measure. A weaker notion is convex risk measure which is a risk measure that satisfies
$\psi(\lambda X+(1-\lambda)Y)\leq\lambda\psi(X)+(1-\lambda)\psi(Y)\quad\forall\lambda\in[0,1]\qquad$ (Convexity) (25)
and finally a risk-measure $\psi$ is called law-invariant if $\psi(X)$ only depends on the distribution of $X$ under $\mathbb{P}$.
We now mention some example of risk measures: The risk measure given by
$\text{ERM}_{\beta}(X)=\frac{1}{\beta}\log\big(\mathbb{E}[e^{-\beta X}]\big)$ (26)
is known as the entropic risk measure with parameter $\beta\neq 0$. It is not positive homogeneous and hence not coherent. For $\beta<0$ it is not even convex. Letting $\beta\rightarrow 0$ one recovers the expectation and letting $\beta\rightarrow\infty$ one recovers the essential infimum risk measure.
The risk measure given by
$\text{VaR}_{\alpha}(X):=q_{\alpha}(X):=\inf\{x\in\mathbb{R}:F_{X}(x)\geq\alpha\}$ (27)
is called the value-at-risk at level $\alpha\in(0,1)$ and is in general not sub-additive hence also not coherent.
The risk measure given by
$\text{CVaR}_{\alpha}(X):=\frac{1}{1-\alpha}\int_{\alpha}^{1}\text{VaR}_{u}(X)du$ (28)
is known as the conditional Value-at-risk (CVaR) or sometimes as the expected shortfall (ES) and is known to be a coherent risk-measure.
All the examples so far are evidently law-invariant.
The actual functional that will be used to rank random variables is the negative of the ERM-risk measure given in the example above with the interpretation being that a lower quantity of risk is preferable. It follows directly then that the functional $\rho:\mathcal{M}\rightarrow\mathbb{R}$ given by $\rho(X):=-\text{ERM}_{\beta}(X)$ has the following properties:
$\rho(0)=0\qquad$ (Normalization) (29)
$\text{if }X\leq Y\text{ then }\rho(X)\leq\rho(Y)\qquad$ (Monotonicity) (30)
$\rho(X+c)=\rho(X)+c\qquad$ (Translation invariance) (31)
It is common in the literature to overload notation and also refer to $\rho$ as the ERM and we will do so and henceforth we will no longer care about risk measures, but only about this specific functional $\rho$. It follows immediately from the normalization and translation invariance that for any real number $c\in\mathbb{R}$ it holds that $\rho(c)=c$.
Appendix C Bellman Optimality and Bellman Recursions
In this section we properly define the state-value functions and state-action value functions of any possibly history-dependent policy $\pi$ and show that the problem of finding an optimal policy can be achieved by a stationary policy and that the value functions satisfy Bellman recursions when the value functions are defined iteratively with respect to the ERM. Several similar results exist in the literature e.g. [52] and [53] which also covers the $\beta>0$ case. These results are derived under more general assumptions on $\mathcal{S}$ and $\mathcal{A}$. These general assumptions are trivially satisfied when $\mathcal{S}$ and $\mathcal{A}$ are finite but their proofs require assumptions on the functionals to ensure the existence of a stationary optimal policy usually by invoking a measurable selection theorem. We avoid this complication by only considering finite $S$ and $A$ and we in turn also give the first proof for state-action value functions and not just for value-functions which is needed as we consider the problem of learning.
Let $M=(\mathcal{S},\mathcal{A},P,R,\gamma,\rho)$ be a finite MDP with $\rho$ being the ERM, and $R(s,a)\in[0,1]$ a deterministic reward function. Let $D=\mathcal{S}\times\mathcal{A}$, $H_{1}=\mathcal{S}$ and $H_{k}=D^{k-1}\times\mathcal{S}$ for $k\geq 2$ the set of all possible histories up to stage $k$. A policy $\pi=(\pi_{k})_{k\in\mathbb{N}}$ a sequence of maps $\pi_{k}:H_{k}\rightarrow\mathcal{A}$. We denote the set of all policies $\Pi$ and identify the set of all stationary policies with the set of measurable maps $F$ from $\mathcal{S}$ to $\mathcal{A}$ which is simply the set of all maps from $\mathcal{S}$ to $\mathcal{A}$ since all maps between finite sets that are both equipped with the discrete topology are measurable with respect to the induced Borel $\sigma$-algebras. Let $B(H_{k})$ be the set of maps $V_{k}:H_{k}\rightarrow\mathbb{R}$ equipped with the supremum norm and let $\pi=(\pi_{k})_{k\in\mathbb{N}}$ be any policy. For any $V_{k+1}\in B(H_{k+1})$ and $h_{k}\in H_{k}$ we denote by $\rho_{h_{k},\pi_{k}}(V_{k+1})$ the functional $\rho$ applied to the random variable concentrated on the set $\{V_{k+1}(h_{k},\pi_{k}(h_{k}),s^{\prime})\}_{s^{\prime}\in S}$ with $\mathbb{P}(s_{k+1}=s^{\prime})=\mathbb{P}(s^{\prime}|s_{k},\pi_{k}(h_{k}))$. By monotonicity of $\rho$ we get that $\rho_{h_{k},\pi_{k}}(V_{k+1})\leq\|V_{k+1}\|$
$(L_{\pi_{k}}V_{k+1})(h_{k})=L_{\pi_{k},V_{k+1}}(h_{k}):=R(s_{k},\pi_{k}(h_{k}))+\gamma\rho_{h_{k},\pi_{k}}(V_{k+1})$ (32)
and similarly we define $L_{a}:B(H_{k+1})\rightarrow B(H_{k})$ by
$(L_{a}V_{k+1})(h_{k})=L_{a,V_{k+1}}(h_{k}):=R(s_{k},a)+\gamma\rho_{s_{k},a}(V_{k+1})$
with $\rho_{s_{k},a}$ defined analogously as $\rho_{h_{k},\pi_{k}}$ as above. By the basic properties of risk-measures it follows directly that $0\leq L_{\pi_{k}V_{k+1}}(h_{k})\leq 1+\gamma\|V_{k+1}\|$ and similarly for $L_{a}$.
For any initial state $s_{1}=s$ we define the $N$-step discounted utility as
$J_{N}(s,a,\pi):=(L_{a}\circ L_{\pi_{2}}\circ\dots\circ L_{\pi_{N}})\boldsymbol{0}(s)$ (33)
where $\boldsymbol{0}(h_{k})=0$ for all $h_{k}\in H_{k}$ and all $k\in\mathbb{N}$.
By monotonicity of $\rho$, it holds that the sequence $(J_{N}(s,a,\pi))_{N\in\mathbb{N}}$ is non decreasing and bounded in the interval $0,\frac{1}{1-\gamma}]$ for any $s,a,\pi\in\mathcal{S}\times\mathcal{A}\times\Pi$ and so the limit
$J(s,a,\pi):=\lim_{N\rightarrow\infty}J_{N}(s,a,\pi)$
exists for any state $s$, any action $a$ and any policy $\pi$.
The problem of the agent is to find $J^{*}(s,a)=\sup_{\pi\in\Pi}J(s,a,\pi)$ and an optimal policy $\pi^{*}$ that solves $J(s,a,\pi^{*})=J^{*}(s,a)$.
Theorem 6.
There exist a unique non-negative function $Q\in B(\mathcal{S}\times\mathcal{A})$ (non-negative maps from $\mathcal{S}\times\mathcal{A}\rightarrow\infty$ equipped with sup-norm) and a stationary decision rule $f^{*}:\mathcal{S}\rightarrow\mathcal{A}$ such that
$Q(s,a)$ $=R(s,a)+\gamma\rho_{s,a}(\max_{a^{\prime}}Q(s^{\prime},a^{\prime}))\,$, (34)
$=R(s,a)+\gamma\rho_{s,a}(Q(s^{\prime},f^{*}(s^{\prime})))\,.$ (35)

### v3

@@ Acknowledgments
Appendix A Risk Measures
In this section, we give a very brief introduction to risk measures. In the actuarial and mathematical finance literature, working with both losses and reward is common. A good introduction is [28], which like us uses the rewards formulation. Due to this ambiguity in the literature, we here collect some precise definitions for the reward setting, and then list some of the most important risk measures.
To begin with, let $(\Omega,\mathcal{F},\mathbb{P})$ be a background probability space, and $\mathcal{M}$ some convex cone of random variables defined on the background space. That is, for any $X,Y\in\mathcal{M}$ and $\lambda>0$, it holds that $X+Y\in\mathcal{M}$ and $\lambda X\in\mathcal{M}$.
Definition 2 (Risk measure).
A functional $\psi:\mathcal{M\rightarrow\mathbb{R}}$ is said to be a risk measure if it satisfies the following properties:
$\psi(0)=0,\qquad$ (Normalization)
$\text{if }X\leq Y\text{ then }\psi(X)\geq\psi(Y),\qquad$ (Monotonicity)
$\psi(X+c)=\psi(X)-c,\quad\forall c\in\mathbb{R}.\qquad$ (Translation invariance)
If, in addition, $\psi$ satisfies the properties
$\psi(cX)=c\psi(X),\quad\forall c>0,\qquad$ (Positive homogeneity)
$\psi(X+Y)\leq\psi(X)+\psi(Y),\qquad$ (Sub-additivity)
it is called a coherent risk measure. A weaker notion is convex risk measure, which is one obeying
$\psi(\lambda X+(1-\lambda)Y)\leq\lambda\psi(X)+(1-\lambda)\psi(Y),\quad\forall\lambda\in[0,1]\,.\qquad$ (Convexity)
Finally, a risk-measure $\psi$ is called law-invariant if $\psi(X)$ only depends on the distribution of $X$ under $\mathbb{P}$.
We now mention some examples of risk measures.
Entropic Risk Measure (ERM).
The risk measure given by
$\text{ERM}_{\beta}(X)=\frac{1}{\beta}\log\big(\mathbb{E}[e^{-\beta X}]\big)$
is known as the entropic risk measure (ERM) with parameter $\beta\neq 0$. Notably, ERM is not coherent (see, e.g., [28]) as it is does not satisfy the positive homogeneity property. Letting $\beta\rightarrow 0$ one recovers the expectation $\mathbb{E}[X]$, and letting $\beta\rightarrow\infty$ yields the essential infimum risk measure.
Value-at-Risk (VaR).
The risk measure given by
$\text{VaR}_{\alpha}(X):=q_{\alpha}(X):=\inf\{x\in\mathbb{R}:F_{X}(x)\geq\alpha\}$
is called the Value-at-Risk (VaR) at level $\alpha\in(0,1)$. VaR is in general not sub-additive, and hence also not coherent.
Conditional Value-at-Risk (CVaR).
The risk measure given by
$\text{CVaR}_{\alpha}(X):=\frac{1}{\alpha}\int_{0}^{\alpha}\text{VaR}_{u}(X)du$
is known as the Conditional Value-at-Risk (CVaR), or sometimes as the expected shortfall (ES). It is known to be a coherent risk-measure.
All the examples so far are evidently law-invariant.
It is worth highlighting that the actual functional $\rho$ used to rank random variables is the negative of the ERM-risk measure $\text{ERM}_{\beta}(X)$, introduced above, with the interpretation being that a lower quantity of risk is preferable. More formally, we consider the functional $\rho:\mathcal{M}\rightarrow\mathbb{R}$ given by $\rho(X):=-\text{ERM}_{\beta}(X)$, featuring the following properties:
$\rho(0)=0,\qquad$ (Normalization)
$\text{if }X\leq Y\text{ then }\rho(X)\leq\rho(Y),\qquad$ (Monotonicity)
$\rho(X+c)=\rho(X)+c.\qquad$ (Translation invariance)
It is common in the literature to overload notation and also refer to $\rho$ as the ERM and we will do so and henceforth we will no longer work directly with risk measures, but only with this specific functional $\rho$. It follows immediately from the normalization and translation invariance that $\rho(c)=c$ for any $c\in\mathbb{R}$.
Appendix B Bellman Optimality and Bellman Recursions
In this section, we properly define the state-value functions and state-action value functions of any possibly history-dependent policy $\pi$ and show that the problem of finding an optimal policy can be achieved by a stationary policy and that the value functions satisfy Bellman recursions when the value functions are defined iteratively with respect to the ERM. Several similar results exist in the literature, e.g., [4] and [6] that also cover the case of $\beta>0$. These results are derived under more general assumptions on $\mathcal{S}$ and $\mathcal{A}$. These general assumptions are trivially satisfied when $\mathcal{S}$ and $\mathcal{A}$ are finite but their proofs require assumptions on the functionals to ensure the existence of a stationary optimal policy usually by invoking a measurable selection theorem. We avoid this complication by only considering finite $\mathcal{S}$ and $\mathcal{A}$ and we in turn also give the first proof for state-action value functions and not just for value-functions, which is needed as we consider the problem of learning.
Let $M=(\mathcal{S},\mathcal{A},P,R,\gamma,\rho)$ be a finite MDP with $\rho$ being the ERM, and $R(s,a)\in[0,1]$ a deterministic reward function. Let $D=\mathcal{S}\times\mathcal{A}$, $H_{1}=\mathcal{S}$ and $H_{k}=D^{k-1}\times\mathcal{S}$ for $k\geq 2$ the set of all possible histories up to stage $k$. A policy $\pi=(\pi_{k})_{k\in\mathbb{N}}$ is a sequence of maps $\pi_{k}:H_{k}\rightarrow\mathcal{A}$. We denote the set of all policies by $\Pi$, and identify the set of all stationary policies with the set of measurable maps $F$ from $\mathcal{S}$ to $\mathcal{A}$, which is simply the set of all maps from $\mathcal{S}$ to $\mathcal{A}$ since all maps between finite sets that are both equipped with the discrete topology are measurable with respect to the induced Borel $\sigma$-algebras. Let $B(H_{k})$ be the set of maps $V_{k}:H_{k}\rightarrow\mathbb{R}$ equipped with the supremum norm, and let $\pi=(\pi_{k})_{k\in\mathbb{N}}$ be any policy. For any $V_{k+1}\in B(H_{k+1})$ and $h_{k}\in H_{k}$, we denote by $\rho_{h_{k},\pi_{k}}(V_{k+1})$ the functional $\rho$ applied to the random variable concentrated on the set $\{V_{k+1}(h_{k},\pi_{k}(h_{k}),s^{\prime})\}_{s^{\prime}\in S}$ with $\mathbb{P}(s_{k+1}=s^{\prime})=\mathbb{P}(s^{\prime}|s_{k},\pi_{k}(h_{k}))$. By monotonicity of $\rho$, we get that $\rho_{h_{k},\pi_{k}}(V_{k+1})\leq\|V_{k+1}\|$.
$(L_{\pi_{k}}V_{k+1})(h_{k})=L_{\pi_{k},V_{k+1}}(h_{k}):=R(s_{k},\pi_{k}(h_{k}))+\gamma\rho_{h_{k},\pi_{k}}(V_{k+1}).$
Similarly, we define $L_{a}:B(H_{k+1})\rightarrow B(H_{k})$ by
$(L_{a}V_{k+1})(h_{k})=L_{a,V_{k+1}}(h_{k}):=R(s_{k},a)+\gamma\rho_{s_{k},a}(V_{k+1}),$
with $\rho_{s_{k},a}$ defined analogously as $\rho_{h_{k},\pi_{k}}$ as above. By the basic properties of risk-measures, it follows directly that $0\leq L_{\pi_{k}V_{k+1}}(h_{k})\leq 1+\gamma\|V_{k+1}\|$ and similarly for $L_{a}$.
For any initial state $s_{0}=s$, we define the $N$-step discounted utility as
$J_{N}(s,a,\pi):=(L_{a}\circ L_{\pi_{2}}\circ\dots\circ L_{\pi_{N}})\boldsymbol{0}(s)$
where $\boldsymbol{0}(h_{k})=0$ for all $h_{k}\in H_{k}$ and all $k\in\mathbb{N}$.
By monotonicity of $\rho$, it holds that the sequence $(J_{N}(s,a,\pi))_{N\in\mathbb{N}}$ is non-decreasing and bounded in the interval $0,\frac{1}{1-\gamma}]$ for any $s,a,\pi\in\mathcal{S}\times\mathcal{A}\times\Pi$, and so the limit
$J(s,a,\pi):=\lim_{N\rightarrow\infty}J_{N}(s,a,\pi)$
exists for any state $s$, any action $a$, and any policy $\pi$.
The agent wishes to find $J^{*}(s,a)=\sup_{\pi\in\Pi}J(s,a,\pi)$ and an optimal policy $\pi^{*}$ attaining $J^{*}(s,a)$, namely, $J(s,a,\pi^{*})=J^{*}(s,a)$.
Theorem 5.
There exist a unique non-negative function $Q\in B(\mathcal{S}\times\mathcal{A})$ (non-negative map from $\mathcal{S}\times\mathcal{A}\rightarrow\infty$ equipped with sup-norm) and a stationary decision rule $f^{*}:\mathcal{S}\rightarrow\mathcal{A}$ such that
$Q(s,a)$ $=R(s,a)+\gamma\rho_{s,a}(\max_{a^{\prime}}Q(s^{\prime},a^{\prime}))\,$,
$=R(s,a)+\gamma\rho_{s,a}(Q(s^{\prime},f^{*}(s^{\prime})))\,.$
Proof.
$LQ(s,a):=R(s,a)+\gamma\rho_{s,a}(\max_{a^{\prime}}Q(s^{\prime},a^{\prime}))\,.$
$LQ_{1}(s,a)$ $-LQ_{2}(s,a)=\gamma[\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime}))-\rho_{s,a}(\max_{a^{\prime}}Q_{2}(s^{\prime},a^{\prime}))]$
$=\gamma[\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime}))-\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime})-\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime})+\max_{a^{\prime}}Q_{2}(s^{\prime},a^{\prime}))]$
$\leq\gamma[\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime}))-\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime})+\max_{a^{\prime}}\{Q_{2}(s^{\prime},a^{\prime})-Q_{1}(s^{\prime},a^{\prime}))\})]$
$\leq\gamma[\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime}))-\rho_{s,a}(\max_{a^{\prime}}Q_{1}(s^{\prime},a^{\prime})+\|Q_{1}-Q_{2}\|)]$
$=\gamma\|Q_{1}-Q_{2}\|\,.$
We start by showing that $L:B(\mathcal{S}\times\mathcal{A})\rightarrow B(\mathcal{S}\times\mathcal{A})$, that is, it takes non-negative functions and returns non-negative functions. By normalization and monotonicity of $\rho$, we have for any non-negative $Q\in B(\mathcal{S}\times\mathcal{A})$ that
$LQ(s,a)=R(s,a)+\gamma\rho_{s,a}(\max_{a^{\prime}}Q(s,a))\geq 0+\gamma\rho_{s,a}(0)=0.$
[…]
Since there are only finitely many states and actions, we can pick a stationary decision rule where $f^{*}(s)$ is an arbitrary element of $\arg\!\max_{a}Q(s,a)$.
Let $V$ be the function given by $V(s):=Q(s,f^{*}(s))$ for all $s$. We then observe that
$V(s)$ $\geq R(s,a)+\gamma\rho_{s,a}(V(s^{\prime}))$
$Q(s,a)\geq(L_{a}\circ L_{\pi_{2}}\circ\cdots\circ L_{\pi_{N}})V(s)\geq(L_{a}\circ L_{\pi_{2}}\circ\cdots\circ L_{\pi_{N}})\boldsymbol{0}(s)=J_{N}(s,a,\pi)\,$,
Finally, we aim to show that $Q(s,a)\leq J(s,a,f^{*})$. By induction, we wish to show that $V(s)\leq J_{N}(s,f^{*}(s),f^{*})+\gamma^{N}\|V\|$ for all $N\in\mathbb{N}$. For the induction step, we start by noting that $J_{1}(s,f^{*}(s),f^{*})=R(s,f^{*}(s))$ and so
$V(s)$ $=R(s,f^{*}(s))+\gamma\rho_{s,f^{*}(s)}(V(s^{\prime}))$
$\leq R(s,f^{*}(s))+\gamma\rho_{s,f^{*}(s)}(\|V\|)$
$=R(s,f^{*}(s))+\gamma\|V\|$
$=J_{1}(s,f^{*}(s),f^{*})+\gamma\|V\|\,$,
$V(s)$ $=L_{f^{*}}V(s)$
$\leq L_{f^{*}}(J_{N}(s,f^{*}(s),f^{*})+\gamma^{N}\|V\|)$
$=\big(R(\cdot,f^{*}(\cdot))+\gamma\rho_{\cdot,f^{*}(\cdot)}(J_{N}(\cdot,f^{*}(\cdot),f^{*})+\gamma^{N}\|V\|\big)(s)$
$=J_{N+1}(s,f^{*}(s),f^{*})+\gamma^{N+1}\|V\|\,$,
from which taking the limit as $N\rightarrow\infty$, we get that $V(s)\leq J(s,f^{*}(s),f^{*})$.
Finally, since
$Q(s,a)=L_{a}V(s)\leq L_{a}J(s,f^{*}(s),f^{*})=J(s,a,f^{*})\,$,
the conclusion holds.
Since this shows that an optimal stationary policy exists, it will suffice to consider only stationary policies and one can by completely analogous arguments show that for any stationary policy $\pi$, there exists a non-negative map $Q^{\pi}\in B(\mathcal{S}\times\mathcal{A})$ such that $Q^{\pi}(s,a)=J(s,a,\pi)$, that is, $Q^{\pi}$ satisfies the Bellman recursion:
$Q^{\pi}(s,a)=R(s,a)+\gamma\rho_{s,a}(Q^{\pi}(s^{\prime},\pi(s^{\prime}))\,$,
We also remark that in the proof, we see directly that $Q(s,a)\in[0,\frac{1}{1-\gamma}]$ for all $(s,a)$.
Appendix C Technical Lemmas
@@
$\log(1-x)$ $\geq-x-x^{2}+x^{3},\quad$ $\forall x\in[0,\frac{1}{5}]$
$\log(1-x)$ $\geq-x-2x,\quad$ $\forall x\in[0,\frac{1}{2}]$
$\log(1+x)$ $\geq x-x^{2},\quad$ $\forall x\in[0,\infty)$
$\log(1+x)$ $\geq\frac{x}{2},\quad$ $\forall x\in[0,1]\,.$
@@
We only prove the first claim, as the rest could be proven using the same technique after some elementary calculations.
Let $f(x)=\log(1-x)$ and $g(x)=-x-x^{2}+x^{3}$. It holds that $f(0)=g(0)$, and since we have $f^{\prime}(x)=-\frac{1}{1-x}$ and $g^{\prime}(x)=-1-2x+3x^{2}$, it follows easily that
$f^{\prime}(x)\geq g^{\prime}(x)\Leftrightarrow 0\leq x(1-5x+3x^{2}),$
@@
Define $f(x)=1-(1-x)^{\alpha}-\frac{x\alpha}{2}$. Since $^{\prime\prime}(x)=-\alpha(\alpha-1)(1-x)^{\alpha-2}<0, ff$ is strictly concave. Further, since $f(0)=0$ and $(\frac{1}{\alpha})=\frac{1}{2}(1-\frac{1}{\alpha})^{\alpha}>\frac{1}{2}-\frac{1}{e}>0, ff$ is positive on the interval $0,\frac{1}{\alpha}]$, and the result follows. ∎
Appendix D Analysis of MB-RS-QVI: Missing Proofs
D.1 Proof of Lemma 1